1. Find the distance between the point and given plane.

P (1, 2, 3)

𝝅: 2x + y – 2z – 4 = 0


1
Expert's answer
2021-07-26T08:58:42-0400
"d=\dfrac{|2(1) + 1(2) – 2(3) – 4|}{\sqrt{(2)^2+(1)^2+(-2)^2}}=2"

The distance between the point P(1,2,3) and given plane is 2 units.